# Mixed-Precision Embeddings for Large-Scale Recommendation Models

### v1

@@ 5.1. Experimental Settings
Following best practices (Li et al., 2023; Lyu et al., 2023), we preprocess the datasets as outlined below. For the Avazu dataset, we remove the instance_id field and transform the timestamp field into three new fields: hour, weekday, and is_weekend. For the Criteo dataset, each numeric value $x$ is discretized into $\lfloor\log^{2}(x)\rfloor$ when $x>2$; otherwise, $x$ is set to 1. For all datasets, categorical features that appear only once are replaced with an ”OOV” (out-of-vocabulary) token. Each dataset is then randomly split in the ratio of 8:1:1 to obtain the corresponding training, validation, and test sets.
@@
We compare the proposed MPE with various embedding compression methods, including QR-Trick (Shi et al., 2020), ALPT (Li et al., 2023), LSQ+ (Bhalgat et al., 2020), PEP (Liu et al., 2021), OptFS (Lyu et al., 2023). QR-Trick is a hashing method, using multiple hash functions and complementary embedding tables to reduce the collisions among features. ALPT is a embedding quantization method, specifically a low-precision training method, which compresses both training and inference memory usage. LSQ+ is an advanced QAT method designed to quantize the weights and activations of convolutional neural networks with a specific bit-width. In this paper, we employ LSQ+ to compress embedding tables of DLRMs. PEP is to prune the feature embeddings with a learnable threshold. OptFS is a feature selection method which learns a gate scalar to decide whether to keep a feature. The implementations of all baselines strictly refer to the original papers or codes, which are available in the open source code repository.
@@ 5.2. Overall Performance (RQ1)
In this section, we primarily address two questions: (1) the maximum compression strength of various methods without losing accuracy, and (2) how the accuracy of these methods varies with different compression ratios. Specifically, we regulate the compression ratio for each method by adjusting relevant hyperparameters, such as the bit-width for LSQ+ and ALPT. However, the QR-trick achieves a minimum compression ratio of 2x and cannot provide lossless compression, making it an exception.
@@
Table 3 presents the accuracy and compression ratios of various methods at lossless compression, which demonstrates that the compression efficiency of MPE significantly surpasses that of other baselines. Specifically, MPE achieves approximately 18, 200, and 20 times compression on embedding tables without comprising prediction accuracy on the three datasets, where the average bit widths are 1.8, 0.16, and 1.6, respectively. The QR-trick shares embeddings across features, which diminishes feature discrimination when multiple features use the same representation. Even with just a 2x compression, the QR-trick leads to significant accuracy loss. LSQ+ achieves a minimum bit-width of 6 without compromising accuracy on all three datasets. Due to the lack of full-precision parameter support during training, ALPT requires a minimum bit-width of 8, which is slightly higher than that of LSQ+. From the pruning perspective, PEP and OptFS prune the embedding table at the element-level and row-level, respectively. They have some compression under the Criteo dataset, while the compression under Avazu and KDD12 datasets is almost negligible, suggesting that there are few useless features or redundant parameters under the latter two datasets. However, LSQ+ can still compress embeddings into 6-bit integer parameters under the Avazu and KDD12 datasets, which indicates that the redundancy in the bitwidth of parameters is significantly higher than that in the number of parameters. In contrast, on the Criteo dataset, the compression performance of LSQ+ is inferior to OptFS and PEP, primarily due to the uniform bit-width applied by LSQ+. Even when some features could be represented with fewer bits, they must use the same bit-width as more important features. Compared to LSQ+, MPE achieves finer-grained compression through the proposed mixed-precision mechanism, assigning different bit-widths to different features. Compared to OptFS and PEP, MPE can also distinguish the importance of different features, and achieves more efficient compression by reducing the precision of embedding parameters. Figure 3 illustrates the accuracy of various methods across different compression ratios, employing the DNN model. Obviously, at the same compression ratio, MPE consistently outperforms all baseline methods in terms of the AUC.
@@ 5.3. Ablation on Retraining (RQ2)
In this section, we examine the impact of retraining on MPE. We evaluated two variants: one without retraining and another with retraining using the LTH, where all parameters are reset to their initial values from the search phase. As shown in Table 4, quantizing embeddings with the sampled bit-widths without retraining leads to a significant loss of accuracy. However, retraining with LTH can partially mitigate this loss, although some degradation in accuracy remains. In MPE, we use the best-performing parameters, except for embeddings, to improve the initialization of the retrained model, resulting in substantial gains in accuracy. Moreover, the impact of retraining varies across datasets, with minimal effects on Criteo and more significant effects on the Avazu and KDD12 datasets.
@@ 5.4. Transferability (RQ3)
In this section, we analyze the transferability of MPE, which specifically refers to the performance of applying the sampled bit-widths from a specific model to other models for retraining. It is important to note that MPE utilizes the step size, offset, and the feature interaction network from the search phase to initialize the model during retraining. However, the structure of feature interaction networks may differ across models. Thus, we only reuse overlapping parameters. The models involved in the search and retraining processes are referred as the source and target models, respectively.
Figure 4 presents the test AUC of the target model using various source models. To evaluate transferability, comparisons should be made within each column, ensuring that the target model remains consistent. On the Criteo dataset, sampled bit-widths can be transferred between different model architectures without impacting retraining accuracy. This aligns with the results in Table 4, which indicate that retraining has little impact on the Criteo dataset. In contrast, for the Avazu and KDD12 datasets, transferring search results across architectures occasionally reduces model accuracy. This implies that feature precision under these datasets is more sensitive to the structure of the feature interaction network. Moreover, the reduction in accuracy can be partially attributed to the inability to fully reuse the parameters of the feature interaction network generated during the search phase. However, compared to the accuracy loss observed without retraining, the loss incurred from transferring search results between different models is negligible.
@@ 5.5. Inference Latency (RQ4)
In this section, we evaluate the inference latency of different methods using the DNN model. The batch size used for inference is 10000. The primary difference among the methods lies in the embedding layer, and the time consumed by subsequent computations remain consistent. Therefore, we represent the inference latency as two parts: embedding table lookup and computation. As illustrated in Figure 5, the time spent on table lookup is significantly shorter than that on data loading and computation, resulting in no substantial difference in overall inference latency across methods. Notably, PEP requires sparse matrix storage for embedding compression, which adds latency to the table lookup process. OptFS reduces the number of features, especially for the Criteo dataset, thereby accelerating table lookup. Methods such as ALPT, LSQ+, and MPE reduce the storage overhead of embedding vectors, also speeding up table lookup. However, the dequantization process of low-precision data slightly offsets the acceleration effect.
@@ 6.1. Embedding Compression
Low-precision methods aim to reduce the precision of embedding parameters through quantization techniques. However, as discussed in Section 1, embedding quantization has received relatively less attention, with most research focusing on low-precision training (LPT) due to its effectiveness in memory compression during training (Xu et al., 2021; Yang et al., 2020; Li et al., 2023). Mixed-dimension methods aim to reduce the size of embedding vectors. For example, CpRec (Sun et al., 2020) assigns smaller embedding dimensions to less frequent features, while MDE (Ginart et al., 2021) adjusts the embedding dimensions according to the size of the associated feature fields. Recent research has utilized neural architecture search (NAS) techniques (Zoph and Le, 2017) to search feature dimensions. For example, ESAPN (Liu et al., 2020) maintains a policy network, optimized via reinforcement learning algorithms, to decide when to increase embedding dimensions. Further, AutoEmb (Zhao et al., 2021a) designs dimension search algorithm through differential architecture search (DARTS) techniques (Liu et al., 2019). Additionally, embedding pruning is essentially a form of mixed-dimension methods. PEP (Liu et al., 2021) prunes embedding vectors by learning a threshold that dynamically determines the embedding size for different features. Weight-sharing methods aim to reduce the number of parameters actually used by embedding tables. For example, DoubleHash (Zhang et al., 2020) allows less frequent features to share embeddings. ROBE applies more advanced hash functions to map each embedding parameter into a shared memory space, enhancing compression efficiency. On the other hand, Wu et al. (2020) employ vector quantization to capture the similarity among features. They first clusters the most frequent embeddings to generate a codebook composed of several codewords. Then, each feature embedding is approximately represented by its most similar codeword. It is important to note that vector quantization is essentially the clustering of vectors, which differs from the parameter quantization discussed in this paper.
@@ 6.2. Mixed-Precision Quantization
In mixed-precision neural networks, prior research has primarily focused on assigning different precision levels across layers to reduce memory consumption and computational overhead (Rakka et al., 2022; Wu et al., 2018; Wang et al., 2019; Dong et al., 2019; Yao et al., 2021). They can be categorized into three categories: gradient-based, heuristic-based, and reinforcement learning-based methods. Gradient-based methods usually convert the discrete precision selection problem into a continuous one, facilitating optimization through gradient descent algorithms. Wu et al. (2018) propose differentiable neural architecture search (DNAS) strategy, which utilizes Gumbel Softmax (Jang et al., 2017) to achieve differentiable precision selection. Subsequent studies, such as BP-NAS (Yu et al., 2020), GMPQ (Wang et al., 2021), and SEAM (Tang et al., 2023), have improved upon DNAS. Heuristic-based methods develop reasonable sensitivity metrics to assign precision. For example, HAWQ (Dong et al., 2019; Yao et al., 2021) uses Hessian information to guide precision assignment, while Hybrid-Net (Chakraborty et al., 2020) applies principal component analysis (PCA) to identify significant layers and assign higher precision accordingly. Reinforcement learning-based methods, like HAQ (Wang et al., 2019) and ADRL (Ning et al., 2021), employ reinforcement learning algorithms to search for optimal precision levels. In this paper, we reformulate the precision search as a precision probability distribution learning problem, where the distribution is updated directly via gradient descent, making it a gradient-based method.
@@ 7. Conclusion
In this paper, we identify a critical drawback of quantization-aware training for embedding compression, that is its inability to distinguish feature importance. To this end, we propose mixed-precision embeddings (MPE) algorithm to enhance the compression efficiency, which identifies an appropriate precision for each feature to balance model accuracy and memory usage. Specifically, we first group features by frequency to simplify the search space for precision, and then learn a probability distribution over precision levels for each feature group. MPE will sample the final bit-widths based on the optimized probability distribution and then retrain the model to achieve better accuracy. Extensive experiments are conducted to evaluate the performance of MPE. The results demonstrate that the MPE significantly outperforms the state-of-the-art methods, achieving a 200-fold compression without any loss in prediction accuracy on the Criteo dataset. Additionally, MPE is implemented as a plug-in embedding layer module based on PyTorch, ensuring ease of use.

### v2

@@ 5.1. Experimental Settings
Following previous work (Li et al., 2023; Lyu et al., 2023), we preprocess the datasets as outlined below. For the Avazu dataset, we remove the instance_id field and transform the timestamp field into three new fields: hour, weekday, and is_weekend. For the Criteo dataset, each numeric value $x$ is discretized into $\lfloor\log^{2}(x)\rfloor$ when $x>2$; otherwise, $x$ is set to 1. For all datasets, categorical features that appear only once are replaced with an ”OOV” (out-of-vocabulary) token. Each dataset is then randomly split in the ratio of 8:1:1 to obtain the corresponding training, validation, and test sets.
@@
We compare the proposed MPE with various embedding compression methods, including QR-Trick (Shi et al., 2020), ALPT (Li et al., 2023), LSQ+ (Bhalgat et al., 2020), PEP (Liu et al., 2021), OptFS (Lyu et al., 2023). QR-Trick is a hashing method, using multiple hash functions and complementary embedding tables to reduce the collisions among features. ALPT is an embedding quantization method, specifically a low-precision training method, which compresses both training and inference memory usage. LSQ+ is an advanced QAT method designed to quantize the weights and activations of convolutional neural networks with a specific bit-width. In this paper, we employ LSQ+ to compress embedding tables of DLRMs. PEP is to prune the feature embeddings with a learnable threshold. OptFS is a feature selection method that learns a gate scalar to decide whether to keep a feature. The implementations of all baselines strictly refer to the original papers or codes, which are available in the open source code repository.
@@ 5.2. Overall Performance (RQ1)
In this section, we primarily address two questions: (1) the maximum compression strength of various methods without losing accuracy, and (2) how the accuracy of these methods varies with different compression ratios. Specifically, we regulate the compression ratio for each method by adjusting relevant hyper-parameters, such as the bit-width for LSQ+ and ALPT. However, the QR-trick achieves a minimum compression ratio of 2x and cannot provide lossless compression, making it an exception.
@@
[…]
AUC under varying compression ratios.
Table 3 presents the accuracy and compression ratios of various methods at lossless compression, which demonstrates that the compression efficiency of MPE significantly surpasses that of other baselines. Specifically, MPE achieves approximately 18, 200, and 20 times compression on embedding tables without comprising prediction accuracy on the three datasets, where the average bit widths are 1.8, 0.16, and 1.6, respectively. The QR-trick shares embeddings across features, which diminishes feature discrimination when multiple features use the same representation. Even with just a 2x compression, the QR-trick leads to significant accuracy loss. LSQ+ achieves a minimum bit-width of 6 without compromising accuracy on all three datasets. Due to the lack of full-precision parameter support during training, ALPT requires a minimum bit-width of 8, which is slightly higher than that of LSQ+. From the pruning perspective, PEP and OptFS prune the embedding table at the element-level and row-level, respectively. They have some compression under the Criteo dataset, while the compression under the Avazu and KDD12 datasets is almost negligible, suggesting that there are few useless features or redundant parameters under the latter two datasets. However, LSQ+ can still compress embeddings into 6-bit integer parameters under the Avazu and KDD12 datasets, which indicates that the redundancy in the bit-width of parameters is significantly higher than that in the number of parameters. In contrast, on the Criteo dataset, the compression performance of LSQ+ is inferior to OptFS and PEP, primarily due to the uniform bit-width applied by LSQ+. Even when some features could be represented with fewer bits, they must use the same bit-width as more important features. Compared to LSQ+, MPE achieves finer-grained compression through the proposed mixed-precision mechanism, assigning different bit-widths to different features. Compared to OptFS and PEP, MPE can also distinguish the importance of different features and achieves more efficient compression by reducing the precision of embedding parameters. Figure 3(c) illustrates the accuracy of various methods across different compression ratios, employing the DNN model. Obviously, at the same compression ratio, MPE consistently outperforms all baseline methods in terms of the AUC.
@@ 5.3. Ablation on Retraining (RQ2)
In this section, we examine the impact of retraining on MPE. We evaluated two variants: one without retraining and another with retraining using the LTH, where all parameters are reset to their initial values from the search phase. As shown in Table 4, quantizing embeddings with the sampled bit-widths without retraining leads to a significant loss of accuracy. However, retraining with LTH can partially mitigate this loss, although some degradation in accuracy remains. In MPE, we use the best-performing parameters, except for embeddings, to improve the initialization of the retrained model, resulting in substantial gains in accuracy. Moreover, the impact of retraining varies across datasets, with minimal effects on the Criteo dataset and more significant effects on the other two datasets.
@@
[…]
Inference latency of the DNN model using different compression methods.
@@ 5.4. Transferability (RQ3)
In this section, we analyze the transferability of MPE, which specifically refers to the performance of applying the sampled bit-widths from a specific model to other models for retraining. It is important to note that MPE utilizes the step size, offset, and feature interaction network from the search phase to initialize the model during retraining. However, the structure of feature interaction networks may differ across models. Thus, we only reuse overlapping parameters. The models involved in the search and retraining processes are referred to as the source and target models, respectively.
Figure 4(c) presents the test AUC of the target model using various source models. To evaluate transferability, comparisons should be made within each column, ensuring that the target model remains consistent. On the Criteo dataset, sampled bit-widths can be transferred between different model architectures without impacting retraining accuracy. This aligns with the results in Table 4, which indicate that retraining has little impact on the Criteo dataset. In contrast, for the Avazu and KDD12 datasets, transferring search results across architectures occasionally reduces model accuracy. This implies that feature precision under these datasets is more sensitive to the structure of the feature interaction network. Moreover, the reduction in accuracy can be partially attributed to the inability to fully reuse the parameters of the feature interaction network generated during the search phase. However, compared to the accuracy loss observed without retraining, the loss incurred from transferring search results between different models is negligible.
@@ 5.5. Inference Latency (RQ4)
In this section, we evaluate the inference latency of different methods using the DNN model. The batch size used for inference is 10000. The primary difference among the methods lies in the embedding layer, and the time consumed by subsequent computations remain consistent. Therefore, we represent the inference latency as two parts: embedding table lookup and computation. As illustrated in Figure 5(c), the time spent on table lookup is significantly shorter than that on data loading and computation, resulting in no substantial difference in overall inference latency across methods. Notably, PEP requires sparse matrix storage for embedding compression, which adds latency to the table lookup process. OptFS reduces the number of features, especially for the Criteo dataset, thereby accelerating table lookup. Methods such as ALPT, LSQ+, and MPE reduce the storage overhead of embedding vectors, speeding up table lookup. However, the dequantization process of low-precision data slightly offsets the acceleration effect.
@@ 6.1. Embedding Compression
Low-precision methods aim to reduce the precision of embedding parameters through quantization techniques. However, as discussed in Section 1, embedding quantization has received relatively less attention, with most research focusing on low-precision training (LPT) due to its effectiveness in memory compression during training (Xu et al., 2021; Yang et al., 2020; Li et al., 2023). Mixed-dimension methods aim to reduce the size of embedding vectors. For example, CpRec (Sun et al., 2020) assigns smaller embedding dimensions to less frequent features, while MDE (Ginart et al., 2021) adjusts the embedding dimensions according to the size of the associated feature fields. Recent research has utilized neural architecture search (NAS) techniques (Zoph and Le, 2017) to search feature dimensions. For example, ESAPN (Liu et al., 2020) maintains a policy network optimized via reinforcement learning algorithms to decide when to increase embedding dimensions. Further, AutoEmb (Zhao et al., 2021a) designs dimension search algorithm through differential architecture search (DARTS) techniques (Liu et al., 2019). Additionally, embedding pruning is essentially a mixed-dimension method. PEP (Liu et al., 2021) prunes embedding vectors by learning a threshold that dynamically determines the embedding size for different features. Weight-sharing methods aim to reduce the number of parameters actually used by embedding tables. For example, DoubleHash (Zhang et al., 2020) allows less frequent features to share embeddings. ROBE applies more advanced hash functions to map each embedding parameter into a shared memory space, enhancing compression efficiency. On the other hand, Wu et al. (2020) employs vector quantization to capture the similarity among features. They first cluster the most frequent embeddings to generate a codebook composed of several codewords. Then, each feature embedding is approximately represented by its most similar codeword. Note that vector quantization is essentially the clustering of vectors, which differs from the parameter quantization discussed in this paper.
@@ 6.2. Mixed-Precision Quantization
In mixed-precision neural networks, prior research has primarily focused on assigning different precision levels across layers to reduce memory consumption and computational overhead (Rakka et al., 2022; Wu et al., 2018; Wang et al., 2019; Dong et al., 2019; Yao et al., 2021). They can be categorized into three categories: gradient-based, heuristic-based, and reinforcement learning-based methods. Gradient-based methods usually convert the discrete precision selection problem into a continuous one, facilitating optimization through gradient descent algorithms. Wu et al. (2018) propose a differentiable neural architecture search (DNAS) strategy, which utilizes Gumbel Softmax (Jang et al., 2017) to achieve differentiable precision selection. Subsequent studies, such as BP-NAS (Yu et al., 2020), GMPQ (Wang et al., 2021b), and SEAM (Tang et al., 2023), have improved upon DNAS. Heuristic-based methods develop reasonable sensitivity metrics to assign precision. For example, HAWQ (Dong et al., 2019; Yao et al., 2021) uses Hessian information to guide precision assignment, while Hybrid-Net (Chakraborty et al., 2020) applies principal component analysis (PCA) to identify significant layers and assign higher precision accordingly. Reinforcement learning-based methods, like HAQ (Wang et al., 2019) and ADRL (Ning et al., 2021), employ reinforcement learning algorithms to search for optimal precision levels. In this paper, we reformulate the precision search as a precision probability distribution learning problem, where the distribution is updated directly via gradient descent, making it a gradient-based method.
@@ 7. Conclusion
In this paper, we identify a critical drawback of quantization-aware training for embedding compression, that is its inability to distinguish feature importance. To this end, we propose the mixed-precision embeddings (MPE) algorithm to enhance the compression efficiency, which identifies an appropriate precision for each feature to balance model accuracy and memory usage. Specifically, we first group features by frequency to simplify the search space for precision and then learn a probability distribution over precision levels for each group. MPE will sample the final bit-widths based on the optimized probability distribution and then retrain the model to achieve better accuracy. Extensive experiments are conducted to evaluate the performance of MPE. The results demonstrate that the MPE significantly outperforms the state-of-the-art methods, achieving a 200-fold compression without any loss in prediction accuracy on the Criteo dataset. Additionally, MPE is implemented as a plug-in embedding layer module based on PyTorch, ensuring ease of use.